OpenStudy (96mertcank96):

In rectangle abcd diagonals ac and bd intersect at point e . if ae= 18 and bd = 2x+30 find x ?

OpenStudy (anonymous):

diagonals of a rectangle are equal. ac=bd ac=2 ae=2*18=36 2x+30=36 find x